OpenStudy (anonymous):

what is the measure of the vertex of an isosceles triangle if one of the base angles measures 58 degrees

OpenStudy (whpalmer4):

Hint: what do you know about the sum of the interior angles in a triangle? How many of the angles in this triangle do you know?

OpenStudy (anonymous):

let the vertex angle be x then x+58+58=180 x=180-116 x=64